OpenStudy (anonymous):

Solve 14^(x+1) = 36 x=____? Round answer to the nearest ten thousandth.

OpenStudy (johnweldon1993):

Hint* \(\large lna^x = x \times ln(a)\) So here...take the log of both sides \[\large (x + 1) \times \ln(14) = \ln(36)\] \[\large x + 1 = \frac{\ln(36)}{\ln(14)}\] \[\large x = \frac{\ln(36)}{\ln(14)} - 1\]

OpenStudy (anonymous):

so answer would be 0.3571?

OpenStudy (johnweldon1993):

I get 0.3579...

OpenStudy (anonymous):

thanks man

OpenStudy (johnweldon1993):

No problem!

OpenStudy (anonymous):

rounded it wrong

OpenStudy (anonymous):

i may have a few more coming up